Use the function to answer the question.
f(x) = 10x-5
What is the value of f –1(–4)?

Answer by stanbon(75887)

Use the function to answer the question.
f(x) = 10x-5
What is the value of f –1(–4)?
----
Find the value of x which makes y = -4
----
Solve 10x-5 = -4
10x = 1
x = 1/10
-----
Therefore f^(-1)(-4) = 1/10
---------------------------
Cheers,
Stan H.